# Machine learning-driven conservative-to-primitive conversion in hybrid piecewise polytropic and tabulated equations of state

### v2

@@ 2.1.1 Piecewise Polytropic EOS Based Model Data
2.1.2 Tabulated EOS Based Model Data
To generate the training data for the tabulated EOS-based model, we sample from a provided EOS table and follow a procedure similar to the one described in Section 2.1.1. We begin by reading in the EOS table, which contains the variables electron fraction ($Y_{e}$), temperature ($T$), rest-mass density ($\rho$), specific internal energy ($\epsilon$), and pressure ($p$). These quantities are stored in logarithmic form in the table and are extracted accordingly. For each data point, a random one-dimensioal three-velocity, $v_{x}$, is sampled from the interval $(0,0.721)$, and values are randomly chosen for electron fraction, temperature, and density from among the tabulated values. Using these, the corresponding values of $\rho$, $\epsilon$, and $p$ are then fetched from the EOS table. The primitive variables are then converted into conserved variables using standard relativistic hydrodynamics relations given in Eq. 2. A total of 1,000,000 datapoints are generated using this process  [33]. Similar to the hybrid piecewise polytropic EOS-based model, the data is split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance before being used for neural network training.
2.2 Model architecture
2.2.1 Piecewise Polytropic EOS Based Model
For the hybrid piecewise polytropic EOS-based model, we tested two feedforward neural networks of varying complexity to represent the conservative-to-primitive variable transformation. Each network takes as input the three conserved variables $(D,S_{x},\tau)$ (Eq. 2) and outputs the pressure $p$ (Eq. 2), assuming the remaining momentum density components are zero for simplicity. After experimenting with multiple multi-layer perceptron (MLP) architectures, we identified two models that offered a good balance between accuracy, speed, and trainability. The smaller model, NNC2PS, features two hidden layers with 600 and 200 neurons, while the larger model, NNC2PL, contains five hidden layers with 1024, 512, 256, 128, and 64 neurons (Fig 2).
@@ 2.2.1 Piecewise Polytropic EOS Based Model
2.2.2 Tabulated EOS Based Model
For the tabulated EOS-based model, we use a single feedforward neural network, NNC2P_Tabulated, to achieve an inherently equivalent task with minor differences. This model takes as input the log-scaled variables $(\log D,\log S_{x},\log\tau,\log Y_{e})$ and outputs the log-scaled pressure, $\log p$ (Eq. 2), assuming $S_{y}$ and $S_{z}$ are zero for simplicity as before. Using log-scaled inputs and outputs aligns with the format of the tabulated EOS values, which are also stored in logarithmic form to accommodate the typically large values of these physical quantities. This approach reduces the range of feature magnitudes, facilitating more stable learning dynamics and better alignment with the source data.
We explored several MLP architectures, varying in parameters, layers, and training strategies, to identify an optimal design for our task. Among these, a structure resembling NNC2PL, detailed in Section 2.2.1 above, emerged as a robust choice. This architecture effectively balanced capacity and efficiency, enabling accurate learning of log-scaled pressure from tabulated EOS data (Fig. 2).
Figure 2: Architectures of the neural networks used for conservative-to-primitive variable mapping. Top: The NNC2PS network takes conserved variables $D$, $S_{x}$, and $\tau$ as input and outputs the pressure $p$. Bottom: The NNC2P_Tabulated network uses the logarithm of conserved variables $\log D$, $\log S_{x}$, and $\log\tau$, along with the electron fraction $Y_{e}$, as input, outputting the logarithm of pressure $\log p$. The NNC2PL network (not shown) shares a similar architecture to NNC2P_Tabulated but with the same input/output structure as NNC2PS.
2.3 Training approach
@@
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}(\theta)-y_{i})^{2}+q\cdot\sum_{i=1}^{n}\mathrm{ReLU}(-\mathcal{N}^{-1}(\hat{y}_{i}(\theta)))\,,$ (5)
@@
All models were trained using the Adam optimizer with an initial learning rate of $3\times 10^{-4}$. A learning rate scheduler reduced the learning rate by a factor of 0.5 if the validation loss failed to improve for five consecutive epochs. NNC2PS and NNC2PL were trained for 85 epochs, while NNC2P_Tabulated required 250 epochs. For each epoch, the model was set to training mode, and data was loaded in batches of 32 onto the GPU. This batch size was chosen based on experimentation to balance the number of epochs and overall time to convergence. While training with larger batches and multiple GPUs (using PyTorch’s DataParallel module or other approaches) is possible, we found no significant advantage regarding the total time to convergence and ultimately opted for this simpler, more portable approach. For each batch, optimizer gradients were reset before generating predictions, and the loss was computed using respective loss functions. Backpropagation was then performed to update the model parameters.
@@
2.4 Inference Speed Tests
In our inference speed tests, we evaluated two main approaches for efficient deployment: a TorchScript model, and NVIDIA’s TensorRT optimized engines. These tests were conducted to measure and compare inference speed under typical deployment conditions, aiming to take advantage of the A100 GPU on Delta.
2.4.1 TorchScript Deployment
@@ 2.4.2 TensorRT Deployment
  • Configuration and Optimization Profiles: The BuilderConfig was set with a 40 GB workspace memory limit, providing more than enough headroom for dynamic batch sizes while maintaining stable performance. We set up a dynamic optimization profile specifying minimum, optimal, and maximum batch sizes within 10 per cent margin of our typical usage, granting flexibility to handle both smaller and larger input volumes efficiently.
@@
In contrast, the TensorRT-based approach demanded several additional configurations. The model, after being converted into an optimized engine, was loaded using TensorRT’s C++ API. This included the manual pre-loading of input data into GPU memory before execution, and was followed by manual setup of input and output buffers for TensorRT’s executeV2 function, and careful management of CUDA resources. While this setup was more involved, it leveraged hardware-specific optimizations to deliver substantial gains in inference speed.
3 Results
3.1 Accuracy
Table 1 summarizes the accuracy results based on $L_{1}$ and $L_{\infty}$ error metrics for each model variant—NNC2PS, NNC2PL, and NNC2P_Tabulated—including both the unquantized and quantized TensorRT engines built from them.
Table 1: Accuracy Results for All Models
@@ 3.1 Accuracy
The NNC2PS model trained in PyTorch achieves very high accuracy with an $L_{1}$ error of $4.54\times 10^{-7}$ and an $L_{\infty}$ error of $3.44\times 10^{-6}$. When the model is converted to a TensorRT engine, the accuracy remains nearly identical, with an $L_{1}$ error of $4.54\times 10^{-7}$ and an $L_{\infty}$ error of $3.43\times 10^{-6}$, indicating minimal loss in precision due to TensorRT optimization. However, when FP16 quantization is applied, the error rates increase to an $L_{1}$ error of $6.39\times 10^{-7}$ and an $L_{\infty}$ error of $8.98\times 10^{-6}$, revealing an obvious side-effect of reduced precision.
The larger NNC2PL model, rather expectedly, achieves lower $L_{1}$ and $L_{\infty}$ errors than NNC2PS, with an $L_{1}$ error of $2.75\times 10^{-7}$ and an $L_{\infty}$ error of $2.61\times 10^{-6}$. The corresponding TensorRT engine preserves this high level of accuracy, showing only a slight and negligible increase to an $L_{1}$ error of $2.88\times 10^{-7}$ and $L_{\infty}$ error of $2.69\times 10^{-6}$ respectively. The FP16 quantized version, however, sees a notable rise in error metrics, with an $L_{1}$ error of $5.32\times 10^{-7}$ and an $L_{\infty}$ error of $9.84\times 10^{-6}$.
The NNC2P_Tabulated model exhibits an $L_{1}$ error of $8.02\times 10^{-3}$ and an $L_{\infty}$ error of $3.54\times 10^{-1}$. It should be noted that the apparent order-of-magnitude difference in its accuracy does not indicate inferiority or failure of the model; in fact, it is the result of the simple fact that this model is trained on a completely different dataset constructed from the LS220 EOS table to estimate the logarithmic pressure values. The TensorRT engine version also shows only a slight increase in $L_{1}$ error to $8.16\times 10^{-3}$. With FP16 quantization, the $L_{1}$ error rises, again, more noticeably to $1.38\times 10^{-2}$.
@@
Additionally, we examined the relative accuracy of the NNC2P_Tabulated model for parameters $W=1.02$, $1.1$, $1.25$, and $1.4$ with $Y_{e}\approx 0.1$ (See Fig. 3). The relative error, defined as the absolute error divided by the true value for each point in a specific parameter set, was not uniform across the parameter space. Larger relative errors were observed in the lowest density and temperature regions of the EOS table, while slightly smaller errors occurred in the high-temperature regions. This accuracy trend was consistent across all tested Lorentz factor ($W$) values and even more emphasized for the FP16 precision TensorRT engine. The LS220 EOS, as provided by  [32], transitions from detailed treatment at high densities to simplified approximations at lower densities, which may contribute to these disparities. Low-density regions are inherently challenging due to the dominance of thermal effects, non-uniform phase transitions, and the treatment of nuclear matter surfaces, which can exacerbate modeling errors [32, 34]. These characteristics likely explain the reduced accuracy in these regions, where variations in the nuclear matter’s phase state are more pronounced.
3.2 Inference Speed Analysis
@@
The inference performance of various methods was evaluated using a single NVIDIA A100 GPU for neural network models and a single-threaded CPU implementation of the traditional numerical method from the RePrimAnd library. The CPUs used in this study were dual AMD 64 core 2.45 GHz Milan processors on the Delta cluster, which can support up to 128 threads. Each configuration was tested across five dataset sizes, ranging from 25,000 to 1,000,000 data points, with ten inference runs conducted per configuration to ensure result stability and consistency.
@@
The numerical method exhibited linear scaling of inference time with respect to the dataset size. In contrast, both TensorRT and TorchScript models generally maintained relatively stable inference times across the dataset sizes. Notably, the full-precision TensorRT engine for the smaller network, NNC2PS, showed a faster-than-expected processing time at certain intermediate dataset sizes, as observed in Fig. 4(a). This behavior may be attributed to favorable thread block utilization and the kernel selection mechanism of TensorRT for this particular network size. A more detailed profiling study is needed to fully elucidate the underlying cause. The accuracy characteristics of these models remained consistent, as indicated in Table 1.
Figure 4: Ideal scaling comparison of various C2P inversion methods under the assumption of perfect parallelization. (a) Projected inference time as a function of dataset size for a traditional numerical solver (RePrimAnd utilizing 128 CPU threads on a single node of the Delta cluster) and two neural network models (NNC2PS and NNC2PL) using TensorRT (FP32 and FP16 precision) and TorchScript across 8 NVIDIA A100 GPUs. (b) Projected inference speed comparison for a dataset of 8 million points, highlighting the significant scalability and efficiency gains achieved by TensorRT engines, particularly with FP16 optimization. The mixed-precision TensorRT engine for NNC2PS achieves approximately a 25-fold reduction in processing time compared to the numerical method, showcasing the potential for TensorRT-based methods to convincingly outperform traditional numerical solvers at scale.
@@
The numerical method required significantly more time than the neural network-based approaches. On average, the numerical method took 103.8 ms to process 25,000 data points, with runtime scaling almost linearly to 3,490 ms for 1,000,000 data points. In contrast, the neural network models demonstrated substantially faster inference times. Specifically, the mixed-precision TensorRT engine built from NNC2PS required 7.92 ms for 25,000 data points and 8.54 ms for 1,000,000 data points. Its full-precision counterpart exhibited similar performance, with runtimes of 25.17 ms for 25,000 data points and 21.06 ms for 1,000,000 data points. The TorchScript variant showed slower performance, but still maintained sub-linear scaling, with runtimes averaging 72.79 ms for 25,000 points and 101.74 ms for 1,000,000 points.
@@
Figure 4 presents the ideal scaling achieved under the assumption of perfect parallelization, providing a theoretical performance benchmark. This scenario assumes optimal workload distribution, minimal communication overhead, and negligible synchronization delays, representing the upper bound of scalability. For the numerical method, the figure reflects the full computational capacity of a single CPU node on the Delta cluster, utilizing 128 threads. For the neural networks, it represents the use of 8 A100 GPUs within a single GPU node. Under these ideal conditions, the processing time of the numerical method per data point is projected to decrease by a factor of 128, allowing for the processing of 8 million points in approximately 218 ms (Fig. 4(b)). Similarly, all neural network methods are expected to achieve linear inference scaling with similar per-GPU efficiency. Under this scenario, TensorRT-based methods—particularly the mixed-precision engine for NNC2PS—show a 25-fold reduction in processing time for 8 million points compared to the numerical method running at full capacity on the CPU node. Furthermore, the scaling trend strongly favors TensorRT for even larger datasets.
@@
4 Conclusions

### v3

@@ II.1.1 Piecewise Polytropic EOS-Based Model Data
II.1.2 Tabulated EOS-Based Model Data
To generate the training data for the tabulated EOS-based model, we sample from a provided EOS table and follow a procedure similar to the one described in Section II.1.1. We begin by reading in the EOS table, which contains the variables electron fraction ($Y_{e}$), temperature ($T$), rest-mass density ($\rho$), specific internal energy ($\epsilon$), and pressure ($p$). These quantities are stored in logarithmic form in the table and are extracted accordingly. For each data point, a random one-dimensional three-velocity, $v_{x}$, is sampled uniformly on a linear scale from the interval $(0,0.721)$. Values for electron fraction and temperature are also sampled uniformly on a linear scale from their respective ranges in the table. The rest-mass density is chosen by randomly selecting one of the grid points from the table, which are logarithmically spaced. For this study, we fetched the corresponding values of $p$ and $\epsilon$ directly from the table without interpolation to ensure the training data perfectly represents the tabulated EOS. Using these, the corresponding values of $\rho$, $\epsilon$, and $p$ are then fetched from the EOS table. The primitive variables are then converted into conserved variables using standard relativistic hydrodynamics relations given in Equation (II). A total of 1,000,000 data points are generated using this process  [32]. Similarly to the hybrid piecewise polytropic EOS-based model, the data is split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance before being used for neural network training.
II.2 Model Architecture
II.2.1 Piecewise Polytropic EOS-Based Model
For the hybrid piecewise polytropic EOS-based model, we tested two feedforward neural networks of varying complexity to represent the conservative-to-primitive variable transformation. Each network takes as input the three conserved variables $(D,S_{x},\tau)$ (Equation (II)) and outputs the pressure $p$ (Equation (II)), assuming the remaining momentum density components are zero for simplicity. This architecture is designed to effectively learn the hidden symmetries in the relationship between the conserved and primitive variables, approximating the intricate C2P transformation without explicit root-finding. After experimenting with multiple multi-layer perceptron (MLP) architectures, as detailed in Appendix A, we identified two models that offered an optimal balance between accuracy, speed, and trainability. The smaller model, NNC2PS, features two hidden layers with 600 and 200 neurons, while the larger model, NNC2PL, contains five hidden layers with 1024, 512, 256, 128, and 64 neurons (Figure 2).
Figure 2: Architectures of the neural networks used for conservative-to-primitive variable mapping. Top: The NNC2PS network takes conserved variables $D$, $S_{x}$, and $\tau$ as input and outputs the pressure pp. Bottom: The NNC2P_Tabulated network uses the logarithm of conserved variables $\log D$, $\log S_{x}$, and $\log\tau$, along with the electron fraction $Y_{e}$, as input, outputting the logarithm of pressure $\log p$. The NNC2PL network shares an identical architecture with NNC2P_Tabulated, but with the input/output structure of NNC2PS.
@@ II.2.1 Piecewise Polytropic EOS-Based Model
II.2.2 Tabulated EOS-Based Model
For the tabulated EOS-based model, we use a single feedforward neural network, NNC2P_Tabulated, to achieve an inherently equivalent task with minor differences. This model takes as input the log-scaled variables $(\log D,\log S_{x},\log\tau,\log Y_{e})$ and outputs the log-scaled pressure, $\log p$ (Equation (II)), assuming $S_{y}$ and $S_{z}$ are zero for simplicity as before. Using log-scaled inputs and outputs aligns with the format of the tabulated EOS values, which are also stored in logarithmic form to accommodate the typically large values of these physical quantities. This approach reduces the range of feature magnitudes, facilitating more stable learning dynamics and better alignment with the source data.
We explored several MLP architectures, varying in parameters, layers, and training strategies, to identify an optimal design for our task. Among these, an architecture identical to NNC2PL, featuring five hidden layers with 1024, 512, 256, 128, and 64 neurons, respectively, detailed in Section II.2.1 above, emerged as a robust choice. This architecture effectively balanced capacity and efficiency, enabling accurate learning of log-scaled pressure from tabulated EOS data (Figure 2).
II.3 Training Approach
@@
[…]
@@
All models were trained using the Adam optimizer with an initial learning rate of $3\times 10^{-4}$. A learning rate scheduler reduced the learning rate by a factor of 0.5 if the validation loss failed to improve for five consecutive epochs. NNC2PS and NNC2PL were trained for 85 epochs, while NNC2P_Tabulated required 250 epochs. These epoch counts were determined empirically by monitoring the validation loss, with training stopped once the loss had clearly converged. The use of a learning rate scheduler, which reduces the learning rate when the validation loss plateaus, also serves as a form of early stopping. For each epoch, the model was set to training mode, and data was loaded in batches of 32 onto the GPU. This batch size was chosen based on experimentation to balance the number of epochs and overall time to convergence. While training with larger batches and multiple GPUs (using PyTorch’s DataParallel module or other approaches) is possible, we found no significant advantage regarding the total time to convergence and ultimately opted for this simpler, more portable approach. For each batch, optimizer gradients were reset before generating predictions, and the loss was computed using respective loss functions. Backpropagation was then performed to update the model parameters.
@@
II.4 Inference Speed Tests
In our inference speed tests, we evaluated two main approaches for efficient deployment: a TorchScript model and NVIDIA’s TensorRT optimized engines. These tests were conducted to measure and compare inference speed under typical deployment conditions, aiming to take advantage of the A100 GPU on Delta.
II.4.1 TorchScript Deployment
@@ II.4.2 TensorRT Deployment
  • Configuration and Optimization Profiles: The BuilderConfig was set with a 40 GB workspace memory limit, providing more than enough headroom for dynamic batch sizes while maintaining stable performance. We set up a dynamic optimization profile specifying minimum, optimal, and maximum batch sizes within a 10 percent margin of our typical usage, granting flexibility to handle both smaller and larger input volumes efficiently.
@@
To ensure we measure the maximum possible performance for each point in our benchmark, we build a specialized, yet flexible, TensorRT engine for each combination of model and dataset size. The dynamic optimization profile for each of these engines is configured with a tight margin around its target dataset size ($N$), as detailed in Table 1.
Table 1: Dynamic optimization profiles used for building specialized TensorRT engines for each benchmarked dataset size (N). Each profile is configured with a tight margin around its target optimal size.
Target Dataset Size (N) Min Batch Size (0.95 N) Optimal Batch Size (N) Max Batch Size (1.05 N)
25,000 23,750 25,000 26,250
50,000 47,500 50,000 52,500
100,000 95,000 100,000 105,000
500,000 475,000 500,000 525,000
1,000,000 950,000 1,000,000 1,050,000
@@
In contrast, the TensorRT-based approach demanded several additional configurations. The model, after being converted into an optimized engine, was loaded using TensorRT’s C++ API. This included the manual pre-loading of input data into GPU memory before execution and was followed by manual setup of input and output buffers for TensorRT’s executeV2 function and careful management of CUDA resources. While this setup was more involved, it leveraged hardware-specific optimizations to deliver substantial gains in inference speed.
III Results
III.1 Accuracy
We evaluate the model accuracy using two standard metrics for regression problems: the $L_{1}$ error (mean absolute error) and the $L_{\infty}$ error (maximum absolute error), both calculated over the entire test dataset. Table 2 summarizes the accuracy results based on $L_{1}$ and $L_{\infty}$ error metrics for each model variant—NNC2PS, NNC2PL, and NNC2P_Tabulated—including both the unquantized and quantized TensorRT engines built from them.
The NNC2PS model trained in PyTorch achieves very high accuracy with an $L_{1}$ error of $4.54\times 10^{-7}$ and an $L_{\infty}$ error of $3.44\times 10^{-6}$. When the model is converted to a TensorRT engine, the accuracy remains nearly identical, with an $L_{1}$ error of $4.54\times 10^{-7}$ and an $L_{\infty}$ error of $3.43\times 10^{-6}$, indicating minimal loss in precision due to TensorRT optimization. However, when FP16 quantization is applied, the error rates increase to an $L_{1}$ error of $6.39\times 10^{-7}$ and an $L_{\infty}$ error of $8.98\times 10^{-6}$, revealing an obvious side effect of reduced precision. This highlights the classic trade-off between computational performance and numerical precision, a critical consideration for selecting the appropriate model for a given scientific application where the tolerance for numerical error may vary.
Table 2: Accuracy results for all models.
@@ III.1 Accuracy
The larger NNC2PL model, rather expectedly, achieves lower $L_{1}$ and $L_{\infty}$ errors than NNC2PS, with an $L_{1}$ error of $2.75\times 10^{-7}$ and an $L_{\infty}$ error of $2.61\times 10^{-6}$. The corresponding TensorRT engine preserves this high level of accuracy, showing only a slight and negligible increase to an $L_{1}$ error of $2.88\times 10^{-7}$ and $L_{\infty}$ error of $2.69\times 10^{-6}$, respectively. The FP16 quantized version, however, sees a notable rise in error metrics, with an $L_{1}$ error of $5.32\times 10^{-7}$ and an $L_{\infty}$ error of $9.84\times 10^{-6}$.
The NNC2P_Tabulated model exhibits an $L_{1}$ error of $8.02\times 10^{-3}$ and an $L_{\infty}$ error of $3.54\times 10^{-1}$. It is important to clarify that this larger error does not indicate a failure of the ML model but is a direct consequence of the model learning from a completely different dataset constructed from the LS220 EOS table to estimate the logarithmic pressure values. The TensorRT engine version also shows only a slight increase in $L_{1}$ error to $8.16\times 10^{-3}$. With FP16 quantization, the $L_{1}$ error rises again, more noticeably, to $1.38\times 10^{-2}$.
Additionally, we examined the relative accuracy of the NNC2P_Tabulated model for parameters $W=1.02$, $1.1$, $1.25$, and $1.4$ with $Y_{e}\approx 0.1$ (See Figure 3). The relative error, defined as the absolute error divided by the true value for each point in a specific parameter set, was not uniform across the parameter space. Larger relative errors were observed in the lowest density and temperature regions of the EOS table, while slightly smaller errors occurred in the high-temperature regions. This accuracy trend was consistent across all tested Lorentz factor ($W$) values and even more emphasized for the FP16 precision TensorRT engine. The LS220 EOS, as provided by  [31], transitions from detailed treatment at high densities to simplified approximations at lower densities, which may contribute to these disparities. Low-density regions are inherently challenging due to the dominance of thermal effects, non-uniform phase transitions, and the treatment of nuclear matter surfaces, which can exacerbate modeling errors [31, 33]. These characteristics likely explain the reduced accuracy in these regions, where variations in the nuclear matter’s phase state are more pronounced.
@@
III.2 Inference Speed Analysis
@@
The inference performance of various methods was evaluated using a single NVIDIA A100 GPU for neural network models and a single-threaded CPU implementation of the traditional numerical method from the RePrimAnd library. The CPUs used in this study were dual AMD 64-core 2.45 GHz Milan processors on the Delta cluster, which can support up to 128 threads. Each configuration was tested across five dataset sizes, ranging from 25,000 to 1,000,000 data points, with ten inference runs conducted per configuration to ensure result stability and consistency. For the RePrimAnd numerical solver, we set the target accuracy for the relative error in the root-finding algorithm to $10^{-8}$. This is a standard, high-precision value used in production codes. We chose to compare our ML models against this robust baseline rather than tuning the numerical solver’s accuracy to match that of the NNs, ensuring a conservative performance comparison.
@@
The numerical method exhibited linear scaling of inference time with respect to the dataset size. In contrast, both TensorRT and TorchScript models generally maintained relatively stable inference times across the dataset sizes. Notably, the full-precision TensorRT engine for the smaller network, NNC2PS, showed a faster-than-expected processing time at certain intermediate dataset sizes, as observed in Figure 4a. This behavior may be attributed to favorable thread block utilization and the kernel selection mechanism of TensorRT for this particular network size. A more detailed profiling study is needed to fully elucidate the underlying cause. The accuracy characteristics of these models remained consistent, as indicated in Table 2.
@@
The numerical method required significantly more time than the neural network-based approaches. On average, the numerical method took 103.8 ms to process 25,000 data points, with runtime scaling almost linearly to 3490 ms for 1,000,000 data points. In contrast, the neural network models demonstrated substantially faster inference times. Specifically, the mixed-precision TensorRT engine built from NNC2PS required 7.92 ms for 25,000 data points and 8.54 ms for 1,000,000 data points. Its full-precision counterpart exhibited similar performance, with runtimes of 25.17 ms for 25,000 data points and 21.06 ms for 1,000,000 data points. The TorchScript variant showed slower performance but still maintained sub-linear scaling, with runtimes averaging 72.79 ms for 25,000 points and 101.74 ms for 1,000,000 points.
Figure 4: Ideal scaling comparison of various C2P inversion methods under the assumption of perfect parallelization. (a) Projected inference time as a function of dataset size for a traditional numerical solver (RePrimAnd utilizing 128 CPU threads on a single node of the Delta cluster) and two neural network models (NNC2PS and NNC2PL) using TensorRT (FP32 and FP16 precision) and TorchScript across 8 NVIDIA A100 GPUs. (b) Projected inference speed comparison for a dataset of 8 million points, highlighting the significant scalability and efficiency gains achieved by TensorRT engines, particularly with FP16 optimization. The mixed-precision TensorRT engine for NNC2PS achieves approximately a 25-fold reduction in processing time compared to the numerical method, showcasing the potential for TensorRT-based methods to convincingly outperform traditional numerical solvers at scale. The width of the lines in panel (a) represents the standard deviation over ten independent runs, with the wider band for the numerical method indicating higher runtime variability.
@@
Figure 4 presents a theoretical performance benchmark based on ideal scaling under the assumption of perfect parallelization. This scenario assumes optimal workload distribution, minimal communication overhead, and negligible synchronization delays, representing the upper bound of scalability. For the numerical method, the figure reflects the full computational capacity of a single CPU node on the Delta cluster, utilizing 128 threads. For the neural networks, it represents the use of 8 A100 GPUs within a single GPU node. Under these ideal conditions, the processing time of the numerical method per data point is projected to decrease by a factor of 128, allowing for the processing of 8 million points in approximately 218 ms (Figure 4b). Similarly, all neural network methods are expected to achieve linear inference scaling with similar per-GPU efficiency. Under this scenario, TensorRT-based methods—particularly the mixed-precision engine for NNC2PS—show a 25-fold reduction in processing time for 8 million points compared to the numerical method running at full capacity on the CPU node. Furthermore, the scaling trend strongly favors TensorRT for even larger datasets.
@@
It is important to contextualize the comparison between the fully utilized CPU component (128 threads) and the fully utilized GPU component (8 GPUs) of a single compute node. This ’node-to-node’ benchmark is designed to answer the practical question of how to best utilize the co-located and often cost-equivalent hardware resources of a modern heterogeneous compute node. While a formal cost-normalized analysis is complex, this approach compares the optimal-use scenario for each hardware type available to a researcher on a typical allocation. The resulting 25-fold speedup is therefore a combination of the algorithmic shift (from iterative root-finding to direct-mapping) and the architectural advantage of GPUs for the massively parallel workload presented by the neural network.
IV Conclusions